# Heterogeneous Intertemporal Treatment Effects via Dynamic Panel Data Models

### v1

@@ 3.3 An Adjusted Inverse-Propensity-Weighted Estimator
There are two keys to the result. First, the mean of the trend $\delta_{2}(D_{1})$ in the decomposition of $\Delta Y_{2}$ in (12) can be identified and removed in the new estimand $E[\widetilde{\Delta Y_{2}}]$ under Assumption 2. Second, the use of inverse propensity weights allows us to attain a more interpretable causal parameter (ATE) without the weights in Proposition 2.
By Proposition 3, we can estimate the ATE, $E[\tau_{2}(D_{1})]$, using sample analogs. Let $\widehat{M}_{2i}$ and $\widehat{M}_{1i}$ be consistent (e.g., kernel-based, Nadaraya-Watson) estimators for
$M_{2i}\equiv E[\Delta Y_{2}|D_{2}=0,Y^{1}=(Y_{i0},Y_{i1}),D_{1}=D_{i1}]$
and
$M_{1i}\equiv E[D_{2}|Y^{1}=(Y_{i0},Y_{i1}),D_{1}=D_{i1}]$
evaluated at any realized $(Y_{i0},Y_{i1},D_{i1})$. The analog estimator for $\mu_{\tau_{2}}\equiv E[\tau_{2}(D_{1})]$ is:
$\widehat{\mu}_{\tau_{2}}\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}\left(\frac{\Delta Y_{i2}-\widehat{M}_{2i}}{\widehat{M}_{1i}}\right)\text{.}$ (20)
In Appendix B, we show this adjusted IPW estimator $\widehat{\mu}_{\tau_{2}}$ is consistent for $\mu_{\tau_{2}}$.
Robins (2000) introduced an IPW estimator for $E[Y_{t}(d^{t})]$ for any $d^{t}\in\{0,1\}^{t}$ in marginal structural models. Despite being motivated by his original IPW estimator, our adjusted IPW estimator targets a different causal parameter. Specifically, our estimand is $E[\tau_{2}(D_{1})]=E[Y_{2}(D_{1},1)-Y_{2}(D_{1},0)]$, while that of Robins’ IPW estimator is $E[Y_{2}(d_{1},d_{2})]$. Two further distinctions of our estimator are that it is based on the first-differenced outcome $\Delta Y_{2}$ and that it pools all observations in estimation. In addition to pooling in estimation, our target parameter provides a potentially useful summary causal effect in multi-period settings where the number of potential contrasts grows large. That is, $E[Y_{t}(D^{t-1},1)-Y_{t}(D^{t-1},0)]$ can be thought of as a weighted average of the $2^{t-1}$ average effects $E[Y_{t}(d^{t-1},1)-Y_{t}(d^{t-1},0)]$ that are in principle identified under sequential exchangeability.
Alternatively, we can construct a numerically equivalent 2SLS estimator for $E[\tau_{2}(D_{1})]$ after transforming $\widetilde{\Delta Y_{2}}$ using the weights $W(Y_{0},D_{1})$. To see this, let
$\Delta Y_{2}^{\dagger}\equiv\widetilde{\Delta Y_{2}}\times\frac{E[W^{2}]}{W},$
and run a transformed 2SLS regression on $\Delta Y_{2}^{\dagger}$. First, regress the endogenous regressors in (8), $\Delta Y_{1}$ and $\Delta D_{2}$, on a fully saturated set of instruments $(Y_{0},D_{1})$ to obtain the fitted values. Then regress $\Delta Y_{2}^{\dagger}$ on the fitted values (and an intercept). It is clear from (13) in Lemma 1 that, provided the projection errors $W(Y_{0},D_{1})$ are nonzero almost surely, such an estimand is equal to $E[\widetilde{\Delta Y_{2}}]$, and therefore $E[\tau_{2}(D_{1})]$ by Proposition 3. In practice, we can implement this transformed 2SLS by first estimating $\Delta Y_{2i}^{\dagger}$ as the product of each individual summand in $\widehat{\mu}_{\tau_{2}}$ and its estimated weights $\frac{1}{n}\sum_{j}\widehat{W}_{j}^{2}/\widehat{W}_{i}$, where $\widehat{W}_{i}$ is the sample analog estimator for the projection errors $w(Y_{0i},D_{1i})$.
3.4 Settings Without Sequential Exchangeability
Sequential exchangeability (Assumption 2) has played a central role in the preceding results of Subsections 3.2 and 3.3. It seems natural to ask whether this assumption is necessary for causal inference in our setting. Indeed, the identification of linear dynamic panel data models with persistent unobserved heterogeneity as in (1) suggests it is not necessary per se, at least if replaced with other (e.g., parametric) assumptions. The goal of this subsection is to establish this in our potential outcome framework and show that for a class of dynamic data-generating processes, standard GMM-based DPDEs recover causal estimands even though sequential exchangeability is generically violated without conditioning on unobservables.
Consider a specialized case in which the series of potential outcomes are structurally determined by a linear dynamic panel data model.
@@ 3.4 Settings Without Sequential Exchangeability
Potential outcomes follow a linear, homogeneous, AR(1) model:
$Y_{it}(d^{t})=\beta^{*}d_{t}+\gamma^{*}Y_{it-1}(d^{t-1})+\theta_{t}^{*}+\alpha_{i}^{*}+\varepsilon_{it}^{*}(d_{t})\text{ for }t\geq 1,$ (21)
with initial condition $Y_{i0}(d^{0})\equiv Y_{i0}$, and where $\beta^{*},\gamma^{*}$, and $\theta_{t}^{*}$ are constant coefficients, $\alpha_{i}^{*}$ are individual-specific fixed effects, and $\varepsilon_{it}^{*}(d_{t})$ are time-varying noises in potential outcomes.
Unlike $\gamma$ in the dynamic panel data model for observed outcomes in (7), the coefficient $\gamma^{*}$ in this model for potential outcomes in (21) has an inherently causal interpretation: it measures how changes in present potential outcomes depend on counterfactual changes in lagged (past) potential outcomes, which depend recursively on the full counterfactual treatment history $d^{t-1}$. Note that the structure of Assumption 5 also implies the potential outcome parametrization of Assumption 1.
The following observation connects the linear dynamic panel data model for potential outcomes to the reduced-form model (7) for observed outcomes.
Observation 1
Suppose the data-generating process for potential outcomes satisfies Assumption 5, and the observed outcomes are $Y_{it}=Y_{it}(D_{it})$. Then the observed sequence $(D_{it},Y_{it})$ satisfies (7) with $(\gamma,\beta,\alpha_{i},\theta_{t})=(\gamma^{*},\beta^{*},\alpha_{i}^{*},\theta_{t}^{*})$ and $\varepsilon_{it}=\varepsilon_{it}^{*}(D_{it})$.
This connection is useful because conditions for consistent estimation of the reduced-form, realized outcome model (7) are well-established. We focus on the Arellano-Bond (1991) GMM estimator, which requires the following set of assumptions on the reduced-form errors $\varepsilon_{it}$:444 Condition (24) is implicit in the moment condition in Arellano and Bond (1991) (Eq. (2) in Sec. 2).
$E(\varepsilon_{is}\varepsilon_{it})=0\text{ for all }t\neq s;$ (22)
$E(D_{is}\varepsilon_{it})=0\text{ for all }s\leq t;$ (23)
$E(\Delta\varepsilon_{it}\alpha_{i})=E(\Delta\varepsilon_{it}Y_{i0})=0\text{ for }t\geq 2.$ (24)
In turn, a sufficient condition on the structural errors $\varepsilon_{it}^{*}(\cdot)$ combines sequential exogeneity and serial uncorrelation, conditional on individual fixed effects. For simplicity, we term this strong structural sequential exogeneity. Let $\varepsilon_{i}^{*t}(d^{t})\equiv(\varepsilon_{i1}^{*}(d_{1}),\dots,\varepsilon_{it}^{*}(d_{t}))$ for $t\geq 1$ and $\varepsilon_{i}^{*t}(\cdot)$ the collection of potential error vectors, with $\varepsilon^{*0}_{i}(\cdot)$ a vacuous null vector.
Assumption 6 (Strong Structural Sequential Exogeneity)
The potential error terms defined under Assumption 5 satisfy:
$E[\varepsilon_{it}^{*}(\cdot)|D_{i}^{t},\varepsilon^{*t-1}_{i}(\cdot),Y_{i0},\alpha_{i}^{*}]=0\text{ for all }t\geq 1.$ (25)
Sufficiency for the Arellano-Bond econometric assumptions is formalized in the following proposition:
@@
Under Assumptions 5 and 6, the Arellano-Bond econometric assumptions (22), (23), and (24) are satisfied.
The next proposition shows that, within the linear dynamic model for potential outcomes (Assumption 5), strong structural sequential exogeneity (Assumption 6) implies a “conditional” version of mean sequential exchangeability that controls for the fixed effects.
Proposition 5
Under Assumptions 5 and 6, mean sequential exchangeability holds conditional on the fixed effect:
for all $d^{t}\in\{0,1\}^{t}$, and where $d^{s-1}$ is a subvector of the first $s-1$ terms of $d^{t}$.
A corollary of Proposition 5 is that the mean sequential exchangeability in Assumption 2b would not hold in general, if the conditional mean in (26) is non-degenerate in individual fixed effect $\alpha_{i}^{*}$, and if the distribution of $\alpha_{i}^{*}$ conditional on history $Y_{i}^{s-1},D_{i}^{s}$ is not invariant in the most recent treatment $D_{is}$. In other words, the unconditional mean sequential exchangeability condition on potential outcomes would not hold in the context of the linear, structural model of Assumption 5 if both potential outcomes and contemporary treatments depend on individual fixed effects in non-trivial ways.
Summarizing, under Assumptions 5 and 6, standard GMM-based DPDEs recover causal estimands (Proposition 4) even though sequential exchangeability is generically violated without conditioning on unobservables (Proposition 5). It is important to reiterate that our results in Section 3.2 and 3.3 do not require such a specific structural model of potential outcomes or econometric restrictions on the errors.

### v2

@@ 3.2 Causal Interpretation of First-Differenced IV Estimands
Proposition 3 establishes the sample average of $\widetilde{\Delta Y_{2}}$ as a standalone estimator of the average treatment effect $\tau_{2}(D_{1})$ under the sequential exchangeability condition. A corollary of Lemma 1 and Proposition 3 is that an adjusted first-differenced IV estimand which replaces $\Delta Y_{2}$ with:
$\Delta Y_{2}^{\dagger}\equiv\widetilde{\Delta Y_{2}}\times\frac{E[w(Y_{0},D_{1})^{2}]}{w(Y_{0},D_{1})}$
also identifies $E[\tau_{2}(D_{1})]$. Unlike Proposition 2, this result holds even without Assumption 3 limiting heterogeneity.
3.3 Causal Inference Without Sequential Exchangeability
Sequential exchangeability (Assumption 2) played a central role in Section 3.2. Yet this condition does not hold generally in observational settings where selection into realized treatments depends on unobserved heterogeneity correlated with contemporary potential and past observed outcomes. (We provide a detailed discussion on this subject in Section 4.2.)
It is therefore natural to ask whether we can recover causal parameters without sequential exchangeability. In this section, we explore other ways to use dynamic panel data models for causal inference where sequential exchangeability does not hold, and yet certain conditional average treatment effects remain (point) identifiable under auxiliary assumptions.
3.3.1 AR(1) Potential Outcome Model
Recall that the dynamic panel data model for observed outcomes in (1) is identified under the conditions of Arellano and Bond (1991), including sequential exogeneity of $D_{it}$ and the serial uncorrelation of the reduced-form errors $\varepsilon_{it}$.
Our goal in Section 3.3.1 is to provide a set of sufficient assumptions on the distribution of potential outcomes and treatments, so that the implied DGP of observed outcomes conform to the functional form of DPDM in (3) and satisfy the conditions in Arellano and Bond (1991). The assumptions we introduce below include homogeneous treatment effects and (conditionally) serially uncorrelated structural errors in potential outcomes. Under such assumptions, the first-differenced IV estimand in (5) recovers the homogeneous treatment effects, even when sequential exchangeability does not hold.
Consider a data-generating process in which the series of potential outcomes follow an AR(1) model with homogeneous causal effects.
Assumption 5 (ARPO)
Potential outcomes follow an AR(1) model:
$Y_{it}(d^{t})=\beta^{*}d_{t}+\gamma^{*}Y_{it-1}(d^{t-1})+\theta_{t}^{*}+\alpha_{i}^{*}+\varepsilon_{it}^{*}(d^{t})\text{ for }t\geq 1,$ (17)
with initial condition $Y_{i0}(d^{0})\equiv Y_{i0}$, [^8] Let $d^{0}$ denote the null vector so that $Y_{i0}(d^{0})\equiv Y_{i0}$ conforms with the notation $Y_{it}(d^{t})\equiv Y_{it}(d_{1},...,d_{t})$. where $\beta^{*},\gamma^{*}$, and $\theta_{t}^{*}$ are constants, $\alpha_{i}^{*}$ are individual fixed effects, and $\varepsilon_{it}^{*}(d^{t})$ are time- and history-varying noises satisfying:
$E[\varepsilon_{it}^{*}(\cdot)|D_{i}^{t},Y_{i0},\varepsilon^{*t-1}_{i}(\cdot),\alpha_{i}^{*}]=0$ (18)
for all $t\geq 1$, where $\varepsilon_{i}^{*t}(d^{t})\equiv(\varepsilon_{i1}^{*}(d^{1}),\dots,\varepsilon_{it}^{*}(d^{t}))$ for $t\geq 1$ and $\varepsilon_{i}^{*t}(\cdot)$ is the collection of potential error vectors with $\varepsilon^{*0}_{i}(\cdot)$ a vacuous null vector.
Unlike $\gamma$ in the DPDM for observed outcomes in (3), the coefficient $\gamma^{*}$ for potential outcomes in (17) has a built-in causal interpretation: it measures how changes in present potential outcomes depend on counterfactual changes in lagged (past) potential outcomes, which depend recursively on the full counterfactual treatment history $d^{t-1}$.
The condition (18) on the structural errors $\varepsilon_{it}^{*}(\cdot)$ combines sequential exogeneity and serial uncorrelation, conditional on individual fixed effects. Note Assumption 5 is also consistent with the condition on potential outcomes in Assumption 1.
The observation below connects the structural model for potential outcomes in (17) to the reduced-form DPDM for observed outcomes in (3).
Observation 1
Suppose the data-generating process for potential outcomes satisfies Assumption 5, and the observed outcomes are $Y_{it}=Y_{it}(D_{i}^{t})$. Then the observed sequence $(D_{i}^{t},Y_{i}^{t})$ satisfies (3) with $(\gamma,\beta,\theta_{t},\alpha_{i})=(\gamma^{*},\beta^{*},\theta_{t}^{*},\alpha_{i}^{*})$ and $\varepsilon_{it}=\varepsilon_{it}^{*}(D_{i}^{t})$.
Our next proposition shows Assumption 5 implies a “conditional” version of mean sequential exchangeability that controls for the fixed effects.
Proposition 4
Under Assumption 5, mean sequential exchangeability (Assumption 2b) holds conditional on the fixed effect:
for all $d^{t}\in\{0,1\}^{t}$, and where $d^{s-1}$ is a subvector of the first s1s-1 terms of $d^{t}$.
Even under Assumption 5, the conditional mean in (19) could be non-degenerate in individual fixed effect $\alpha_{i}^{*}$, whose distribution conditional on history $Y_{i}^{s-1},D_{i}^{s}$ may depend on the most recent treatment $D_{is}$. In such cases, Assumption 2b does not hold in general, because both potential outcomes and contemporary treatments depend on individual fixed effects in non-trivial ways.
Despite possible violation of sequential exchangeability, Assumption (5) is sufficient for consistent estimation of the structural parameters $(\beta^{*},\gamma^{*},\Delta\theta_{2}^{*})$ using the Arellano-Bond IV estimand in (5), as applied to a reduced-form DPDM for observed outcomes in (3).
Proposition 5
[…]
Generalization to the identification of later trends $\Delta\theta^{*}_{t}$ for $t\geq 3$ is straightforward; it only requires using $X_{it}\equiv(\Delta D_{it},\Delta Y_{it-1},1)$ and $Z_{it}\equiv E(X_{it}|Y_{i}^{t-2},D_{i}^{t-1})$ in the first-differenced IV estimand.
The intuition for Proposition 5 is that (18) implies the standard Arellano-Bond assumptions on reduced-form errors $\varepsilon_{it}$ are satisfied. The result then follows from Observation 1.
In summary, under Assumptions 5, a first-differenced IV estimand recovers the causal estimands (Proposition 5), even when sequential exchangeability is violated without conditioning on unobservables (Proposition 4). Unlike Section 3.2, this result does require a structural model of potential outcomes in (17) as well as econometric restrictions on the errors in potential outcomes.
3.3.2 AR(1) Untreated Potential Outcome Model
In this subsection, we obtain stronger and more robust results for identifying causal effects, generalizing from the homogeneous treatment effects in the ARPO model (Assumption 5). Specifically, we let the treatment effects be heterogeneous, time-varying, and history-dependent:
$\tau_{it}^{*}(d^{t-1})\equiv Y_{it}(d^{t-1},1)-Y_{it}(d^{t-1},0)$ (20)
and impose the structure and sequential exogeneity only on untreated potential outcomes. We refer to this as an AR(1) untreated potential outcomes (ARUPO) model. Besides the heterogeneous treatment effects, we recycle coefficient labeling to avoid introducing new notation.
Assumption 6 (ARUPO)
Untreated potential outcomes follow an AR(1) model:
$Y_{it}(d^{t-1},0)=\gamma^{*}Y_{it-1}(d^{t-1})+\alpha^{*}_{i}+\theta^{*}_{t}+\varepsilon^{*}_{it}(d^{t-1})\text{ for }t\geq 1,$ (21)
with initial condition $Y_{i0}(d^{0})\equiv Y_{i0}$, where $\gamma^{*}$ and $\theta_{t}^{*}$ are constants, $\alpha_{i}^{*}$ are individual fixed effects, and $\varepsilon_{it}^{*}(d^{t-1})$ are time- and history-varying noises in potential outcomes satisfying:
$E[\varepsilon^{*}_{it}(d^{t-1})|D_{i}^{t}=d^{t},Y_{i0},\tau_{i}^{*t-1}(d^{t-2}),\varepsilon_{i}^{*t-1}(\cdot),\alpha^{*}_{i}]=0,$ (22)
for all $t\geq 1$ and $d^{t}\in\{0,1\}^{t}$.
This condition allows $D_{it}$ to be a function of the most recent outcome $Y_{it-1}$ and the full past history $(D_{i}^{t-1},Y_{i}^{t-2})$. It posits the noises in potential outcomes $\varepsilon^{*}_{it}(\cdot)$ are uncorrelated with the sequence of past treatments up to $D_{i}^{t}$ and all latent elements that determine the past observed outcomes up to $Y_{i}^{t-1}$.
The structure of the ARUPO model (Assumption 6) implies the observed outcomes are generated by a linear dynamic panel data model with a random coefficient for contemporary treatment:
$Y_{it}\equiv Y_{it}(D_{i}^{t})=\tau_{it}^{*}D_{it}+\gamma^{*}Y_{it-1}+\alpha_{i}^{*}+\theta_{t}^{*}+\varepsilon_{it}^{*},$
where $\tau_{it}^{*}$ and $\varepsilon_{it}^{*}$ are shorthand for $\tau^{*}_{it}(D_{i}^{t-1})$ and $\varepsilon^{*}_{it}(D_{i}^{t-1})$, respectively. First-differencing the observed outcomes yields:
$\Delta Y_{it}=\tau_{it}^{*}D_{it}-\tau_{it-1}^{*}D_{it-1}+\gamma^{*}\Delta Y_{it-1}+\Delta\theta_{t}^{*}+\Delta\varepsilon_{it}^{*}\text{ for }t\geq 2.$ (23)
Under Assumption 6,
$E(\varepsilon_{it}^{*}|D_{i}^{t},Y_{i}^{t-1})=0\text{ for all }t\geq 1,$
because the observed history $Y_{i}^{t-1}$ is a function of $(D_{i}^{t-1},Y_{i0},\tau_{i}^{*t-1}(D_{i}^{t-2}),\varepsilon_{i}^{*t-1}(D_{i}^{t-2}),\alpha^{*}_{i})$. This implies the observed history of treatment and outcomes is sequentially exogenous:
$E(\Delta\varepsilon_{it}^{*}|H_{it-1})=0,\text{ where }H_{it-1}\equiv(D_{i}^{t-1},Y_{i}^{t-2}).$ (24)
This is reminiscent of the moment condition implied by sequentially exogenous regressors in Arellano and Bond (1991). However, the ARUPO model is a generalization over the DPDM for observed outcomes: the first-difference of observed outcomes in (23) involves random coefficients (heterogeneous treatment effects) $\tau_{it}^{*}D_{it}-\tau_{it-1}^{*}D_{it-1}$ instead of the product of a constant coefficient and $\Delta D_{it}$.
We first focus on identifying the average treatment effect of the first-time-treated defined for $t\geq 1$:
$ATFT_{t}(y_{0})\equiv E[\tau_{it}^{*}|D_{it}=1,\mathcal{H}_{t-1}(y_{0})].$ (25)
where for simplicity of notation we let:
$\mathcal{H}_{t-1}(y_{0})\equiv\{D_{i}^{t-1}=0^{t-1},Y_{i0}=y_{0}\}$
denote the event of an untreated history among those with initial outcome $y_{0}$. In Appendix B.1 we extend our approach to also identify similar causal parameters with a general history of past treatments $d^{t-1}\neq 0^{t-1}$.
To begin, note that if the untreated model parameters $(\gamma^{*},\Delta\theta_{t}^{*})$ could be identified, then the $ATFT_{t}(y_{0})$ would also be identified through (23) and (24) upon taking conditional expectations and simplifying:
$E[\Delta Y_{it}|\mathcal{H}_{it-1}(y_{0})]=E[D_{it}|\mathcal{H}_{it-1}(y_{0})]ATFT_{t}(y_{0})+\gamma^{*}E[\Delta Y_{it-1}|\mathcal{H}_{it-1}(y_{0})]+\Delta\theta_{t}^{*}.$ (26)
Yet, it is not apparent how to identify these untreated model parameters in isolation without further assumptions.
We show how to identify $ATFT_{t}(y_{0})$, using a conditional Arellano-Bond IV estimand for (23) under an additional condition on how the observed treatments relate to latent components in the ARUPO model (Assumption 6), in particular the heterogeneous treatment effects.
@@ 3.3.2 AR(1) Untreated Potential Outcome Model
For any initial condition $y_{0}$,
$E[\tau^{*}_{it}(0^{t-1})|D_{it},\mathcal{H}_{t-1}(y_{0}),\varepsilon_{i}^{*t-1}(\cdot),\alpha^{*}_{i}]=E[\tau^{*}_{it}(0^{t-1})|D_{it},\mathcal{H}_{t-1}(y_{0})].$
Remarkably, this condition does not impose a notion of conditional unconfoundedness, because it does not rule out the correlation between the current treatment $D_{it}$ and the heterogeneous treatment effects $\tau^{*}_{it}(0^{t-1})$, which contribute to the contemporary potential outcome $Y_{it}(0^{t-1},1)$. Instead, it only states the fixed effect $\alpha^{*}_{i}$ and past noises $\varepsilon_{i}^{*t-1}(\cdot)$ do not affect the mean treatment effect conditional on the observed history. Of course, this assumption is also trivially satisfied in the ARPO model (Assumption 5) where the treatment effects are homogeneous.
To define (a family of) conditional IV estimands, let $w_{t-1}(\cdot)$ be a chosen function of $H_{it-1}$ and write $W_{it-1}\equiv w_{t-1}(H_{it-1})$. Next, fix an initial condition value $Y_{0i}=y_{0}$ and consider an IV regression of $\Delta Y_{it}$ on $X_{it}\equiv(\Delta D_{it},\Delta Y_{it-1},1)$ conditional on the event $\mathcal{H}_{t-1}(y_{0})$, using the following instruments:
$Z_{it}(y_{0})\equiv E[X_{it}|W_{it-1},\mathcal{H}_{t-1}(y_{0})].$
These instruments are exogenous, because (24) implies $E[\Delta\varepsilon_{it}^{*}|Z_{it},\mathcal{H}_{t-1}(y_{0})]=0$. The estimand in such a (conditional) IV regression is:
$\begin{pmatrix}\widetilde{\beta}_{t}(y_{0})\\ \widetilde{\gamma}_{t}(y_{0})\\ \widetilde{\Delta\theta_{t}}(y_{0})\end{pmatrix}\equiv E[Z_{it}^{\prime}(y_{0})X_{it}|\mathcal{H}_{t-1}(y_{0})]^{-1}E[Z_{it}^{\prime}(y_{0})\Delta Y_{it}|\mathcal{H}_{t-1}(y_{0})]$ (27)
Our next result shows that under Assumptions 6 and 7, the IV regression in (27) recovers the causal parameter of interest, $ATFT_{t}(y_{0})$, as well as the (constant) state dependence parameter $\gamma^{*}$ and the time trends $\Delta\theta_{t}^{*}$.
@@
Suppose Assumptions 6 and 7 hold, and $E[Z_{it}^{\prime}(y_{0})Z_{it}(y_{0})|\mathcal{H}_{t-1}(y_{0})]$ has full rank. Then the IV estimands (27) for $t\geq 3$ recover the causal terms:
$\widetilde{\beta}_{t}(y_{0})$ $=ATFT_{t}(y_{0});$
$\widetilde{\gamma}_{t}(y_{0})$ $=\gamma^{*};$
$\widetilde{\Delta\theta_{t}}(y_{0})$ $=\Delta\theta_{t}^{*}.$
Both $\gamma^{*}$ and $\Delta\theta_{t}^{*}$ are over-identified since the estimands are functions of $y_{0}$. The intuition for Proposition 6 is as follows. Under Assumption 7, the heterogeneous treatment effects $\tau^{*}_{it}$ are mean-independent from $Y_{i}^{t-2}$ conditional on $D_{it}$ and $\mathcal{H}_{t-1}(y_{0})$. This implies the conditional mean of $\Delta Y_{it}$ given $W_{it-1}$ and $\mathcal{H}_{t-1}(y_{0})$ is linear in the instruments $Z_{it}(y_{0})$, with slope coefficients $ATFT_{t}(y_{0})$, $\gamma^{*}$, and an intercept $\Delta\theta_{t}^{*}$.999 That is, an equality similar to (26) holds after further conditioning on $W_{it-1}$, which is excluded from $ATFT_{t}(y_{0})$ by Assumption 7. As a result, these parameters are recovered via an OLS regression of $\Delta Y_{it}$ on $Z_{it}(y_{0})$ conditional on $\mathcal{H}_{t-1}(y_{0})$. By construct, such an estimand also coincides with that of a conditional IV regression in (27), because $E[Z_{it}^{\prime}(y_{0})X_{it}(y_{0})|\mathcal{H}_{t-1}(y_{0})]=E[Z_{it}^{\prime}(y_{0})Z_{it}(y_{0})|\mathcal{H}_{t-1}(y_{0})]$ by the law of iterated expectations.
In Appendix B.1 we generalize Assumption 7 and Proposition 6 to recover the average treatment effects beyond the first-treated:
$ATT_{t}(d^{t-1},y_{0})\equiv E[\tau^{*}_{it}(d^{t-1})|D_{it}=1,D_{i}^{t-1}=d^{t-1},Y_{i0}=y_{0}]\text{ for }d^{t-1}\neq 0^{t-1}.$
A natural further generalization of the ARUPO model in Assumption 6 would be to replace the constant $\gamma^{*}$ with random coefficients $\gamma_{i}^{*}$, which vary by individual but not over time; in turn, a natural extension of (18) conditions exogeneity on the extended vector of fixed effects. Even so, the analysis of Chamberlain (2022) suggests identification problems in this case of multiple random coefficients, with only partial identification of causal effects possible (e.g., see also Lee (2022)), even for the untreated outcome model in isolation. Therefore we do not consider this case further in this paper.